# Particle identification in the GlueX detector with machine learning

### v2

@@ Abstract
In particle physics experiments, identifying the types of particles registered in a detector is essential for the accurate reconstruction of particle collisions. At Thomas Jefferson National Accelerator Facility (Jefferson Lab), the GlueX experiment performs particle identification (PID) by setting specific thresholds, known as cuts, on the kinematic properties of tracks and showers obtained from detector hits. Our research aims to enhance this cut-based method by employing machine-learning algorithms based on multi-layer perceptrons and boosted decision trees. Similar approaches have been applied in other particle physics experiments and offer an opportunity to increase PID accuracies using reconstructed kinematic data. Our study illustrates that both MLPs and BDTs can identify charged and neutral particles in Monte Carlo (MC) simulated GlueX data with significantly improved accuracy over the current cuts-based PID method.
@@ 1 Introduction
The GlueX experiment, located in Hall D at Jefferson Lab, searches for evidence of gluonic degrees of freedom in the excitation spectrum of light mesons through the mapping of the spectrum of mesons with exotic quantum numbers with masses up to 2.5 $\nicefrac{{\mbox{GeV}}}{{c^{2}}}$. Accurate identification of these mesons requires knowledge of their production mechanisms, quantum numbers ($J^{PC}$), and decay modes. This is accomplished through the partial wave analysis (PWA) of exclusive final states. To accomplish this, the GlueX experiment relies upon high statistics, linear polarization of the incident photon beam, excellent measurement resolution, full acceptance in decay angles, and correct decay particle identification (PID).
@@
  • SC: A start counter consisting of a cylindrical array of 40 scintillator paddles immediately surrounding the liquid hydrogen target
  • TOF: A forward time-of-flight wall of 44 vertical and 44 horizontal scintillator paddles located in the forward direction downstream of the target
  • BCAL: A cylindrical geometry barrel calorimeter with alternating layers of lead and scintillating fiber located at large polar angles surrounding the target
  • FCAL: A forward lead-glass block calorimeter with a planar geometry orthogonal to the photon beam axis, located in the far-forward region downstream of the target
  • CDC: A central drift chamber detector consisting of layers (strung in axial and stereo configurations) of cathode straw tubes, each containing an anode wire and filled with a mixture of argon and CO2 gas subtending polar angles greater than 20 surrounding the target
  • FDC: A forward drift chamber detector consisting of 24 planar drift chambers with cathode strip and wire readouts, and filled with a mixture of argon and CO2 gas subtending polar angles less than 30 downstream of the target
  • DIRC: A Detection of Internally Reflected Cherenkov light detector used for forward region PID, located directly in front of the forward TOF
@@
In this work, we aim to demonstrate the potential of ML [4] as a tool to improve PID accuracies for GlueX analyses. ML has already been applied to PID in other high energy experiments [5, 6, 7, 8, 9] and has even been applied to a GlueX study [10] that used a discriminating neural network to reduce background in the FCAL. Motivated by such studies and the demand for high-accuracy PID, this work makes use of a trained Multi-Level Perceptron (MLP) and Boosted Decision Tree (BDT) to classify single hadronic particles using GlueX MC simulation data. Cuts-based PID methods currently utilized by GlueX are carried out on the same MC simulation data sets and compared to the PID accuracies obtained from the charged particle MLP and BDT models.
@@ 2 Monte Carlo Simulation Data Set
The training and test data sets used for this study were extracted from low-momentum GlueX particle gun simulations (see Table 1). In these simulations, a particle is spawned at a randomized location, known as a vertex, within the 30 cm long liquid H2 target cell. The generated particle is launched from the assigned vertex in a random direction and a random momentum less than 1 $\nicefrac{{\mbox{GeV}}}{{c}}$. Detailed simulations of interactions between the generated particle and the GlueX detector and any subsequent decays and interactions are performed using the standard GlueX GEANT4-based software package [11, 12]. The resulting simulated detector signals (“hits”) are then reconstructed using the halld_\_recon package for shower and track identification. The kinematic parameters describing each of the reconstructed particles are saved for each event, together with the labels that identify the actual particle type and momentum that were fed as input to the simulation for each event. The complete list of features for each event that are provided to the models as input is shown in Table 2.
@@
E GeV Particle total energy -5
px GeV/c Particle momentum X-component -500
py GeV/c Particle momentum Y-component -500
pz GeV/c Particle momentum Z-component -500
@@ E GeV Particle total energy -5
dEdxCDC keV/cm Average dE/dx of track in the CDC -5
dEdxFDC keV/cm Average dE/dx of track in the FDC -5
@@
dEdxTOF keV/cm Average track dE/dx in the TOF -5
@@
dEdxSc keV/cm dE/dx of track in the SC -5
@@ 4.1 Cuts-Based PID
Table 3: The optimized $\nicefrac{{dE}}{{dx}}$ vs $p$ boundary parameters where $f_{3}$ separates electrons from pions/muons, $f_{2}$ separates pions/muons/electrons from kaons, and $f_{1}$ separates kaons from protons/antiprotons.
@@
Particle $\Delta t$ BCAL [ns] $\Delta t$ FCAL [ns] $\nicefrac{{dE}}{{dx}}$ $\nicefrac{{\mbox{keV}}}{{\mbox{cm}}}$] $\nicefrac{{E}}{{p}}$ [$c$]
@@
Each optimized decision boundary is shown overlaid on the test data set in Figure 2. An additional manual cut is made for electrons and muons/pions, in which the ratio of the particle’s total energy E to the magnitude of the momentum is taken, with a decision boundary at $\nicefrac{{E}}{{p}}=0.83\,c$. A comprehensive summary of the cuts-based PID parameters for charged particles is shown in Table 4. A PID is made for every hypothesis in our test data set that meets the $\chi^{2}$ criteria; however, if a given event met none of the criteria, then no ID is designated. Furthermore, if an event has two or more PIDs that match different hypotheses in the test data set, the particle type with the highest $\chi^{2}$ value is designated.
@@ 6 Conclusion
In addition to higher PID accuracies, the ML models are interpreted using SHAP values to study the most important features of the reconstructed events in the identification of specific particle types. Such an analysis gives insight into how the ML models make particle classifications and provides insight into which characteristics of the reconstructed events are most important in distinguishing one particle type from another. One result of interest from the model interpretation shows that for all charged particles, the distance between the vertex to BCAL is more import for ML PID than the average dE/dx of a track inside of the CDC. Such insights from ML interpretation in future studies using real particle data has the potential to deepen understanding of relationships between each particle and their signature in the GlueX detector.

### v3

@@ Abstract
In particle physics experiments, identifying the types of particles registered in a detector is essential for the accurate reconstruction of particle collisions. At Thomas Jefferson National Accelerator Facility (Jefferson Lab), the GlueX experiment performs particle identification (PID) by setting specific thresholds, known as cuts, on the kinematic properties of tracks and showers obtained from detector hits. Our research aims to enhance this cut-based method by employing machine-learning algorithms based on multi-layer perceptrons and boosted decision trees. Similar approaches have been applied in other particle physics experiments and offer an opportunity to increase PID accuracies using reconstructed kinematic data. Our study illustrates that both multilayered perceptrons and boosted decision trees can identify charged and neutral particles in Monte Carlo simulated GlueX data with significantly improved accuracy over the current cuts-based PID method.
@@ 1 Introduction
The GlueX experiment, located in Hall D at Jefferson Lab, searches for evidence of gluonic degrees of freedom in the excitation spectrum of light mesons through the mapping of the spectrum of mesons with exotic quantum numbers with masses up to 2.5 $\mathrm{GeV}/c^{2}$. Accurate identification of these mesons requires knowledge of their production mechanisms, quantum numbers ($J^{PC}$), and decay modes. This is accomplished through the partial wave analysis (PWA) of exclusive final states. To accomplish this, the GlueX experiment relies upon high statistics, linear polarization of the incident photon beam, excellent measurement resolution, full acceptance in decay angles, and correct decay particle identification (PID).
@@
  • SC: a start counter consisting of a cylindrical array of 40 scintillator paddles immediately surrounding the liquid hydrogen target;
  • TOF: a forward time-of-flight wall of 44 vertical and 44 horizontal scintillator paddles located in the forward direction downstream of the target;
  • BCAL: a cylindrical geometry barrel calorimeter with alternating layers of lead and scintillating fiber located at large polar angles surrounding the target;
  • FCAL: a forward lead-glass block calorimeter with a planar geometry orthogonal to the photon beam axis, located in the far-forward region downstream of the target;
  • CDC: a central drift chamber detector consisting of layers (strung in axial and stereo configurations) of cathode straw tubes, each containing an anode wire and filled with a mixture of argon and CO2 gas subtending polar angles greater than 20 surrounding the target;
  • FDC: a forward drift chamber detector consisting of 24 planar drift chambers with cathode strip and wire readouts, and filled with a mixture of argon and CO2 gas subtending polar angles less than 30 downstream of the target;
  • DIRC: a Detection of Internally Reflected Cherenkov light detector used for forward region PID, located directly in front of the forward TOF.
@@
In this work, we aim to demonstrate the potential of ML [4] as a tool to improve PID accuracies for GlueX analyses. ML has already been applied to PID in other high energy experiments [5, 6, 7, 8, 9] and has even been applied to a GlueX study [10] that used a neural network for classification to reduce background in the FCAL. Motivated by such studies and the demand for high-accuracy PID, this work makes use of a trained Multi-Level Perceptron (MLP) and Boosted Decision Tree (BDT) to classify single hadronic particles using GlueX MC simulation data. Cuts-based PID methods currently utilized by GlueX are carried out on the same MC simulation data sets and compared to the PID accuracies obtained from the charged particle MLP and BDT models.
@@ 2 Monte Carlo Simulation Data Set
The training and test data sets used for this study were extracted from low-momentum GlueX particle gun simulations (see Table 1). In these simulations, a particle is spawned at a randomized location, known as a vertex, within the 30 cm long liquid H2 target cell. The generated particle is launched from the assigned vertex in a random direction and a random momentum less than 1 $\mathrm{GeV}/c$. Detailed simulations of interactions between the generated particle and the GlueX detector and any subsequent decays and interactions are performed using the standard GlueX GEANT4-based software package [11, 12]. The resulting simulated detector signals (“hits”) are then reconstructed using the halld_\_recon package for shower and track identification. The kinematic parameters describing each of the reconstructed particles are saved for each event, together with the labels that identify the actual particle type and momentum that were fed as input to the simulation for each event. The complete list of features for each event that are provided to the models as input is shown in Table 2.
@@
E $\mathrm{GeV}$ Particle total energy -5
px $\mathrm{GeV}/c$ Particle momentum X-component -500
py $\mathrm{GeV}/c$ Particle momentum Y-component -500
pz $\mathrm{GeV}/c$ Particle momentum Z-component -500
@@ E $\mathrm{GeV}$ Particle total energy -5
dEdxCDC $\mathrm{keV}/\mathrm{cm}$ Average $\mathrm{d}E/\mathrm{d}x$ of track in the CDC -5
dEdxFDC $\mathrm{keV}/\mathrm{cm}$ Average $\mathrm{d}E/\mathrm{d}x$ of track in the FDC -5
@@
dEdxTOF $\mathrm{keV}/\mathrm{cm}$ Average track $\mathrm{d}E/\mathrm{d}x$ in the TOF -5
@@
dEdxSc $\mathrm{keV}/\mathrm{cm}$ $\mathrm{d}E/\mathrm{d}x$ of track in the SC -5
@@ 4.1 Cuts-Based PID
Table 3: The optimized energy deposition $(\mathrm{d}E/\mathrm{d}x)$ vs $p$ boundary parameters where $f_{3}$ separates electrons from pions/muons, $f_{2}$ separates pions/muons/electrons from kaons, and $f_{1}$ separates kaons from protons/antiprotons.
@@
Particle $\Delta t$ BCAL [ns] $\Delta t$ FCAL [ns] $\mathrm{d}E/\mathrm{d}x$ $\mathrm{keV}/\mathrm{cm}$] $E/p$ [$c$]
@@
Each optimized decision boundary is shown overlaid on the test data set in Figure 2. An additional manual cut is made for electrons and muons/pions, in which the ratio of the particle’s total energy E to the magnitude of the momentum is taken, with a decision boundary at E/p=0.83cE/p=0.83\,c. A comprehensive summary of the cuts-based PID parameters for charged particles is shown in Table 4. A PID is made for every hypothesis in our test data set that meets the $\chi^{2}$ criteria; however, if a given event met none of the criteria, then no ID is designated. Furthermore, if an event has two or more PIDs that match different hypotheses in the test data set, the particle type with the highest $\chi^{2}$ value is designated.
@@ 6 Conclusion
In addition to higher PID accuracies, the ML models are interpreted using SHAP values to study the most important features of the reconstructed events in the identification of specific particle types. Such an analysis gives insight into how the ML models make particle classifications and provides insight into which characteristics of the reconstructed events are most important in distinguishing one particle type from another. One result of interest from the model interpretation shows that for all charged particles, the distance between the vertex to BCAL is more import for ML PID than the energy deposition $(\mathrm{d}E/\mathrm{d}x)$ of a track inside of the CDC. Such insights from ML interpretation in future studies using real particle data has the potential to deepen understanding of relationships between each particle and their signature in the GlueX detector.